OpenStudy (anonymous):

Evaluate the following limit:

OpenStudy (anonymous):

\[\lim_{x \rightarrow \pi/2}\frac{ 5(\csc x-1) }{ \cot ^{2}x}\]

OpenStudy (anonymous):

@mathteacher1729

OpenStudy (anonymous):

use L hopital rule -5/2

OpenStudy (anonymous):

I think you should break the denominator into parts, like: cot^2 = csc ^2 -1 = (csc +1)(csc-1) and then cancel out with the numerator. then it's easier to calculate

OpenStudy (anonymous):

Thanks all :)

OpenStudy (anonymous):

ok. you can handle it then, right?

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

yea its 5/2

OpenStudy (anonymous):

I don't calculate. you do it.

OpenStudy (anonymous):

lol ok thats what i got